OpenStudy (anonymous):

Simplify [(x-4)^2 - (sqrt(5))^2].

OpenStudy (anonymous):

x^2-8x+16-5 which is x^2-8x+11

OpenStudy (anonymous):

Is there any chance you can tell me how you did that?

OpenStudy (anonymous):

(x-4)^2=x^2-8x+16 which is from(a-b)^2=a^2-2ab+b^2 and 5^1/2*2 the twos get canceled leaving u with only 5 thus x^2-8x+16-5 which is x^2-8x+11

OpenStudy (anonymous):

[(x-4)^2 - (sqrt(5))^2] (x-4)(x-4) - 5 x^2-4x-4x+16 -5 x^2-8x+16-5 x^2-8x+11

OpenStudy (anonymous):

if theres still a part you dont understand feel free to ask

OpenStudy (anonymous):

Thank you both so much!